OpenStudy (anonymous):

What are the differences in measuring the molar mass using freezing point depression between studying ionic solids and molecular solids?

OpenStudy (aaronq):

Molecular solids do not ionize into several particles while ionic solids do (to some extent). Since Freezing-point depression is a colligative property, caused by the presence of particles of non-volatile solute (and not on their chemical nature), ionic solids have a multiplicative effect on the change in temperature per unit - i.e. the van't hoff constant, \(i\). FP depression equation: \(\large \Delta T=i*\underbrace{m}_{molality}*K_b\) Molecular solids have \(i\)=1 always, and ionic solids \(i\geq2. \)

OpenStudy (aaronq):

i forgot to mention that you would rearrange the formula to find the molar mass, like this: \(\Delta T=i*m*K_b\) sub in \(m=\dfrac{n_{solute}}{kg_{solvent}}\) \(\Delta T=i*(\dfrac{n_{solute}}{kg_{solvent}})*K_b\) sub in \(n=\dfrac{m}{M}\) \(\Delta T=i*(\dfrac{\dfrac{m}{M}}{kg_{solvent}})*K_b\) Then you get: \(\Delta T=\dfrac{i*m*K_b}{kg_{solvent}*M}\) Rearrange for molar mass, M: \(M=\dfrac{i*m*K_b}{kg_{solvent}*\Delta T}\)

OpenStudy (aaronq):

NOTE the first \(m\) is molality and the second \(m\) is mass.